# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ 4.5 Modular Decomposition
Modular decompositions exist at the other end of the spectrum of decomposability, where individual value functions and/or policies can be learned for each decomposed entity $X$. Specifically, a task can be broken down into individual subsystems $\kappa_{1},\dots,\kappa_{N}$ for which models, value functions, and policies can be subsequently learned. Such modularity can exist along the following axes: (i) Spatial Modularityallows learning quantities specific to parts of the state space, thus, effectively reducing the dimensionality of the states; (ii) Temporal Modularityallows breaking down tasks into sequences over a learning horizon and, thus, learning modular quantities in a sequence; (iii) Functional Modularityallows decomposing the policy architecture into functionally modular parts, even if the problem is spatially and temporally monolithic.
A potential consequence of such breakdown is the emergence of a hierarchy, and when learning problems exploit this hierarchical relationship, then these problems come under the purview of Hierarchical RL (HRL) (?). The learned policies can also exhibit a hierarchy, where each policy can choose the lower-level policies to execute the subtasks. Each level can be treated as a planning problem (?) or a learning problem (?), thus, allowing solutions to combine both planning and learning through the hierarchy. Hierarchy, however, is not a necessity for modularity.
@@
Modular decomposition of state spaces is primarily studied at high-level planning and state-abstractions for HRL methods (?). Additionally, the literature on skills has looked into the direction of training policies for individual parts of the state-space (?). Goals have been specifically considered in methods that either use goals as an interface between levels of hierarchy (?, ?, ?), or as outputs of task specification methods (?, ?).
@@
Modularity in action spaces refers to conditioning policies on learned action abstraction. The classic example of such methods belongs to the realm of the Options framework where policies are associated with temporal abstractions over actions (?). In HRL methods, learning and planning of the higher levels are based on the lower-level policies and termination conditions of their execution.
@@
Continual settings utilize policies compositionally by treating already learned policies as primitives. Such methods either feed these primitives to the discrete optimization problems for selection mechanisms or to continuous optimization settings involving ensembling (?) and distillation (?). Modularity in such settings manifests itself by construction and serves as a central factor in building solutions. Even though the final policy in such paradigms, obtained through ensembling, selection, and/or distillation, can be monolithic, the method of obtaining such policies is a purely distributed regime.
@@ 5 Patterns of Incorporating Structure
Having defined different forms of decomposability and the different objectives that side information can be used to accomplish, we now connect the two by understanding the ways of incorporating structure into a learning process. We assume that some form of structure exists in the problem and/or the solution space, which can be incorporated into the learning pipeline as an inductive bias. To understand how decomposability can be incorporated into the RL pipeline, we survey the literature with a very specific question in mind: Do existing methods use structure in a repeatable manner? The answer to this question, inspired by the categorization of ? (?) for the supervised learning case, brings us to patterns of incorporating structure.
@@
Figure 4: Patterns of incorporating structural information. We categorize the methods of incorporating structure as inductive biases into the learning pipeline into patterns that can be applied for different kinds of usages.
A pattern is a principled change in the RL pipeline $\Omega$ that allows the pipeline to achieve one, or a combination of, the additional objectives: Sample Efficiency, Generalization, Safety, and Interpretability. We categorize the literature into seven patterns, an overview of which has been shown in Figure 4. Our proposed patterns come from our survey of the literature and are meant to provide an initial direction for such categorization. We do not consider this list exhaustive but more as a starting point to build further upon. We present an overview of our meta-analysis on what patterns are used for which of the four use cases in Figure 5.
Figure 5: Proclivities. A meta-analysis of the proclivities of each pattern to the additional objectives. On the x-axis are the patterns discussed in this text, while on the y-axis are the percentage of publications for each additional objective that address it using a particular pattern.
In the following subsections, we delve deeper into each pattern, explaining different lines of literature that apply each pattern for different use cases. To further provide intuition about this categorization, we will consider the running example of a taxi service, where the task of the RL agent (the taxi) is to pick up passengers from various locations and drop them at their desired destinations within a city grid. The agent receives a positive reward when a passenger is successfully dropped off at their destination and incurs a small penalty for each time step to encourage efficiency.
For each of the following sections, we present a table of the surveyed methods that categorizes the work in the following manner: (i) The structured space, information about which is incorporated as side information; (ii) The type of decomposition exhibited for that structured space. We specifically categorize works that use structured task distributions through goals and/or rewards; (iii) The additional objectives for which the decomposition is utilized. In addition to demonstrating our categorization, our rationale behind the table format is to highlight the areas where further research might be lucrative. These are the spots in the tables where we could not yet find literature and/or we believe additional work can be important.
@@ 5.1 Abstraction Pattern
Abstraction pattern utilizes structural information to create abstract entities in the RL pipeline. For any entity, $X$, an abstraction utilizes the structural information to create $X_{abs}$, which takes over the role of $X$ in the learning procedure. In the taxi example, the state space can be abstracted to the current grid cell of the taxi, the destination grid cell of the current passenger, and whether the taxi is currently carrying a passenger. This significantly simplifies the state space compared to representing the full details of the city grid. The action space could also be abstracted to moving in the four cardinal directions, plus picking up and dropping off a passenger. Finding appropriate abstractions can be a challenging task in itself. Too much abstraction can lead to loss of critical information, while too little might not significantly reduce complexity (?). Consequently, learning-based methods that jointly learn abstractions factor this granularity into the learning process.
Abstractions have been thoroughly explored in the literature, with early work addressing a formal theory on state abstractions (?). Recent works have primarily used abstractions for tackling generalization. Thus, we see in Figure 5 that generalization is the most explored use case for abstractions. However, the aforementioned advantages of abstraction usually interleave these approaches with sample efficiency gains and safety as well. Given the widespread use of abstractions in the literature, we explore how different forms of abstractions impact each use case in the following paragraphs.
@@
Goals Latent
Relational

### v3

@@ 5.5 Modular Decomposition
Modular decompositions exist at the other end of the spectrum of decomposability, where individual value functions and/or policies can be learned for each decomposed entity $\mathcal{X}$. Specifically, a task can be broken down into subsystems $\mathcal{X}_{1},\dots,\mathcal{X}_{N}$ for which models, value functions, and policies can be independently learned.
$Z:\mathcal{X}\to\{\mathcal{X}_{1},\dots,\mathcal{X}_{N}\}.$ (11)
Such modularity can exist along the following axes: (i) Spatial Modularityallows learning quantities specific to parts of the state space, thus effectively reducing the dimensionality of the states; (ii) Temporal Modularityallows breaking down tasks into sequences over a learning horizon and, thus, learning modular quantities in a sequence; (iii) Functional Modularityallows decomposing the policy architecture into functionally modular parts, even if the problem is spatially and temporally monolithic.
A potential consequence of such breakdown is the emergence of a hierarchy, and when learning problems exploit this hierarchical relationship, these problems come under the purview of Hierarchical RL (HRL) (?). The learned policies can also exhibit a hierarchy, where each can choose the lower-level policies to execute the subtasks. Each level can be treated as a planning problem (?) or a learning problem (?), thus allowing solutions to combine planning and learning through the hierarchy. Hierarchy, however, is not a necessity for modularity.
@@
Modular decomposition of state spaces is primarily studied at high-level planning and state-abstractions for HRL methods (?). Approaches such as Q-decomposition (?, ?) have explored agent design by communicating Q values learned by individual agents on parts of the state-action space to an arbitrator that suggests the following action. Additionally, the literature on skills has looked into the direction of training policies for individual parts of the state-space (?). Similarly, partial models only make predictions or specific parts of the observation-action spaces in Model-Based settings (?, ?). Goals have been considered explicitly in methods that either use goals as an interface between levels of hierarchy (?, ?, ?), or as outputs of task specification methods (?, ?).
@@
Modularity in action spaces refers to conditioning policies on learned action abstractions. The classic example of such methods belongs to the options framework where policies are associated with temporal abstractions over actions (?). In HRL methods, learning and planning of the higher levels are based on the lower-level policies and termination conditions of their execution.
@@
Continual settings utilize policies compositionally by treating already learned policies as primitives. Such methods feed these primitives to the discrete optimization problems for selection mechanisms or to continuous optimization settings involving ensembling (?) and distillation (?). Modularity in such settings manifests itself by construction and is a central factor in building solutions. Even though the final policy in such paradigms, obtained through ensembling, selection, and/or distillation, can be monolithic, obtaining such policies is a purely distributed regime.
@@ 6 Patterns of Incorporating Structure
Having defined different forms of decomposability and the different objectives that side information can be used to accomplish, we now connect the two by understanding the ways of incorporating structure into a learning process. We assume that some form of structure exists in the problem and/or the solution space, which can be incorporated into the learning pipeline as an inductive bias. To understand how decomposability can be incorporated into the RL pipeline, we could start a potential categorization along two axes: the type of decomposition (Latent, Factored, Relational, and Modular) and the part of the pipeline to which this decomposition is applied (such as states, or actions). However, such a categorization ignores a significant part of the process: the method by which the pipeline is conditioned on side information. For example, information about goals can be used to learn state abstractions or directly fed as input to the policy network. Both of these design decisions can have very different impacts in practice. Thus, in addition to the two axes mentioned above, we survey the literature with another specific question: Do existing methods incorporate structural information using repeatable design decisions? The answer to this question, inspired by the categorization proposed by ? (?), brings us to patterns of incorporating structure.
@@
Figure 4: Patterns of incorporating structural information. We categorize the methods of incorporating structure as inductive biases into the learning pipeline into patterns that can be applied for different kinds of usages. Each pattern is shown as a plug-and-play modification of the RL pipeline $\Omega$ that aims to improve the performance of $\Omega$ on one or more objectives discussed in Section 4.
A pattern is a principled change in the RL pipeline $\Omega$ that allows the pipeline to achieve one, or a combination of, the additional objectives: Sample Efficiency, Generalization, Safety, and Interpretability. We categorize the literature into seven patterns, an overview of which has been shown in Figure 4.
Our proposed patterns come from our literature survey and are meant to provide an initial direction for such categorization. We do not consider this list exhaustive but more as a starting point upon which to build further. We present an overview of our meta-analysis on the patterns used for which of the four use cases in Figure 5.
Figure 5: Proclivities. A meta-analysis of the proclivities of each pattern to the additional objectives. The four additional objectives covered in this text are on the x-axis. We show each objective’s share percentage of publications utilizing individual patterns. This data has been shown on the y-axis with different colors for each pattern. Therefore, this figure helps us understand correlations between patterns and the kind of objectives they address.
In the following subsections, we delve deeper into each pattern, explaining different lines of literature that apply each pattern for different use cases. To further provide intuition about this categorization, we will consider the running example of a taxi service, where the task of the RL agent (the taxi) is to pick up passengers from various locations and drop them at their desired destinations within a city grid. The agent receives a positive reward when a passenger is successfully dropped off at their destination and incurs a minor penalty for each time step to encourage efficiency.
For each of the following sections, we present a table of the surveyed methods that categorizes the work in the following manner: (i) The structured space, information about which is incorporated as side information; (ii) The type of decomposition exhibited for that structured space. We specifically categorize works that use structured task distributions through goals and/or rewards; (iii) The additional objectives for which the decomposition is utilized. Our rationale behind the table format is to highlight the areas where further research might be lucrative in addition to categorizing the existing literature. These are the spots in the tables where we could not yet find literature, and we believe additional work can be important; therefore, in addition to categorizing existing methods, empty areas in the table highlight avenues for future research.
@@ 6.1 Abstraction Pattern
Abstraction pattern utilizes structural information to create abstract entities in the RL pipeline. For any entity, $X$, an abstraction utilizes the structural information to create $X_{abs}$, which takes over the role of $X$ in the learning procedure. In the taxi example, the state space can be abstracted to the taxi’s current grid cell, the destination grid cell of the current passenger, and whether the taxi is currently carrying a passenger. This significantly simplifies the state space compared to representing the full details of the city grid. The action space could also be abstracted to moving in the four cardinal directions, plus picking up and dropping off a passenger. Behavioral abstractions are intricately related to history-based abstractions in that they address similar applications, namely, abstraction states and histories, into lower dimensional representations that can be leveraged for RL. Therefore, we address both of these in the following sections.
Figure 6: Abstraction Pattern
Finding appropriate abstractions can be a challenging task in itself. Too much abstraction can lead to loss of critical information, while too little might not significantly reduce complexity (?). Consequently, learning-based methods that jointly learn abstractions factor this granularity into the learning process.
Abstractions have been thoroughly explored in the literature, with early work addressing a formal theory on state abstractions (?, ?). Recent works have primarily used abstractions for tackling generalization. Thus, we see in Figure 5 that generalization is the most explored use case for abstractions. However, the advantages mentioned earlier of abstraction usually interleave these approaches with sample efficiency gains and safety. Given the widespread use of abstractions in the literature, the following paragraphs explore how different abstractions impact each use case.
@@
Goals Latent ? (?) ? (?), ? (?)
Relational ? (?)
Modular ? (?) ? (?) ? (?), ? (?)
States Latent ? (?), ? (?), ? (?), ? (?), ? (?), ? (?), ? (?), ? (?) ? (?), ? (?), ? (?), ? (?), ? (?), ? (?) ? (?) ? (?), ? (?)
Factored ? (?) ? (?), ? (?), ? (?), ? (?), ? (?), ? (?) ? (?), ? (?), ? (?)
Relational ? (?), ? (?), ? (?), ? (?), ? (?) ? (?), ? (?), ? (?), ? (?), ? (?), ? (?), ? (?) ? (?), ? (?)
Modular ? (?), ? (?), ? (?) ? (?), ? (?), ? (?), ? (?) ? (?), ? (?)
Actions Latent ? (?), ? (?)
Factored ? (?) ? (?)
Relational ? (?) ? (?)
Modular ? (?) ? (?), ? (?) ? (?)
Rewards Latent ? (?), ? (?), ? (?), ? (?)
Factored ? (?) ? (?), ? (?), ? (?), ? (?) ? (?)
Dynamics Latent ? (?) ? (?), ? (?), ? (?), ? (?)
Factored ? (?) ? (?)
Modular ? (?) ? (?)
Survey of literature utilizing the abstraction pattern
Generalization.
State abstractions are a standard choice for improving generalization performance by capturing shared dynamics across MDPs into abstract state spaces using methods such as Invariant Causal Prediction (?, ?), similarity metrics (?, ?, ?, ?, ?, ?, ?), Free Energy Minimization (?), and disentanglement (?, ?, ?, ?, ?).
Value functions serve as temporal abstractions for shared dynamics in Multi-task Settings. Successor Features (SF) (?, ?) exploit latent reward and dynamic decompositions using value functions as an abstraction. Subsequent works have combined them with Generalized Policy Iteration (?) and Universal Value Function Approximators (?, ?). Factorization in value functions, on the other hand, helps improve sample efficiency and generalization both (?, ?).
Relational abstractions contribute to generalization by incorporating symbolic spaces into the RL pipeline. These help incorporate planning approaches in hierarchical frameworks (?, ?). Additionally, relational abstractions can help abstract away general aspects of a collection of MDPs, thus allowing methods to learn generalizable Q-values over abstract states and actions that can be transferred to new tasks (?) or develop methods specifically for graph-structured spaces (?, ?).
Abstractions can additionally enable generalization in hierarchical settings by compressing state spaces (?), abstract automata (?, ?), templates of dynamics across tasks (?), or even be combined with options to preserve optimal values (?).
Sample Efficiency.
Latent variable models improve sample efficiency across the RL pipeline. Latent state abstractions help improve sample efficiency in Model-based RL (?) and also help improve the tractability of policy learning over options in HRL (?). In model-free tasks, these are also learned as inverse models for visual features (?) or control in a latent space (?). Latent transition models demonstrate efficiency gains by capturing task-relevant information in noisy settings (?), by preserving bisimulation distances between original states (?), or by utilizing factorized abstractions (?). Learned latent abstractions (?) also contribute to the exploration mechanism in the Go-Explore regime (?).
Latent action models expedite convergence of policy gradient methods such as REINFORCE (?) by shortening the learning horizon in stochastic scenarios like dialog generation (?). Action embeddings, on the other hand, help reduce the dimensionality of large action spaces (?)
Safety and Interpretability.
Relational abstractions are an excellent choice for interpretability since they capture interactionally complex decompositions. The combination of object-centric representations and learned abstractions adds transparency (?) while symbolic interjections, such as tracking the relational distance between objects, help improve performance  (?).
State and reward abstractions help with safety. Latent states help to learn safe causal inference models by embedding confounders (?). On the other hand, meshes (?, ?) help benchmark metrics such as robustness in a learned policy.
6.2 Augmentation Pattern
The augmentation pattern treats $X$ and $z$ as separate input entities for the action-selection mechanism. The combination can range from the simple concatenation of structural information to the state or actions to more involved methods of conditioning policy or value functions on additional information. Crucially, the structural information neither directly influences the optimization procedure nor changes the nature of $X$. It simply augments the already existing entities. In this view, abstractions learned in an auxiliary manner and concatenated to states, actions, or models can also be considered augmentations since the original entity remains unchanged.
Figure 7: Augmentation Pattern
For the taxi example, one way to apply the augmentation pattern would be by conditioning the policy on additional information, such as the time of day or day of the week. This information could be helpful because traffic conditions and passenger demands can vary depending on these factors. However, augmentations can increase the complexity of the policy and the learning process, and care needs to be taken to ensure that the policy does not overfit the additional information. Consequently, this pattern is generally not explored as much as abstraction. While we see usages of augmentations equitably for most use cases in Figure 5, the number of papers utilizing this pattern still falls short compared to more established techniques, such as abstraction. In the following paragraphs, we delineate three augmentations in the surveyed work.
Context-based Augmentations.
Contextual representations of dynamics (?, ?) and goal-related information (?, ?) help with generalization and sample efficiency by exposing the agent to the necessary information for optimality. Goal augmentations additionally allow interpretable mechanisms for specifying goals (?). On the other hand, augmentation of meta-learned latent spaces to the normal state can promote temporally coherent exploration across tasks (?). Action histories (?) can directly help with sample efficiency, and action relations (?, ?) contribute to generalization over large action sets.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Latent ? (?), ? (?)
Factored ? (?) ? (?)
Relational ? (?) ? (?), ? (?)
Modular ? (?), ? (?) ? (?), ? (?) ? (?)
States Latent ? (?), ? (?), ? (?) ? (?), ? (?), ? (?)
Factored ? (?)
Relational ? (?) ? (?)